Assume exports increase by $500 million and imports decrease by $300 million and that the MPC is .75. What's the effect on real GDP?

Answer Choices:
a. $3.2 billion increase
b. $800 million increase
c. $200 million decrease
d. $4 billion increase
1
Expert's answer
2016-03-31T08:57:04-0400
b. $800 million increase
GDP increasing=(Export-Import)*multiplicator of expenses.
multiplicator of expenses=1/(1-MPC)
GDP increasing=(500-300)/(1-0.75)=800